Given the function f described by f(x)=12x, find each of the following.
(a) f(-6)
(b) f(6)
(c) f(21)
f(-6)=___?

Found 2 solutions by solver91311, ewatrrr:
Answer by solver91311(24713)



Take whatever is inside the parentheses and put it in place of , then do the arithmetic.



So







See?

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by ewatrrr(24785)

 
Hi
f(x)=12x
(a) f(-6) = 12·-6 = -72 ||Idea is to simply replace x with value given
(b) f(6) = 12·6 = 72
(c) f(21) = = 12·21 = 252